OpenStudy (anonymous):

Solve this system of equations: 5x+4y=-8 and 2x-y=-11

OpenStudy (anonymous):

method or answer?

OpenStudy (anonymous):

Method and answer

OpenStudy (anonymous):

or maybe the answer, then show work

OpenStudy (anonymous):

multiply second equation by 4, add to first equation

OpenStudy (amistre64):

y = (za-tc)/(ad-bc)

OpenStudy (anonymous):

\[5x+4y=-8\] \[8x-4y=-44\] \[13x=-52\] \[x=-4\]

OpenStudy (amistre64):

5 4 =-8 2 -1 =-11 y = (5(-11)-2(-8))/(5(-1)-2(4))

OpenStudy (anonymous):

then replace x by -4 in either equation, solve for y

OpenStudy (amistre64):

i get y = 3 if i did it right :)

OpenStudy (anonymous):

5x+4y=-8 2x-y=-11 get the bottom to have some common system like below: 5x+4y=-8 8x-4y=-44 the 4's cancel out then add going down: 13x = -52 divide both sides by 13 to get x. x = -4 now plug in -4 in to the first equation 5(-4)+4y=-8 -20+4y=-8 +20 +20 4y=12 divide both sides by 4. you get: y = 3 So your answer is: x = -4, y = 3

OpenStudy (amistre64):

4 5 =-8 -1 2 =-11 x = (4(-11)+1(-8))/(4(2)-5(-1)) = -4 yay... it works